If four consecutive odd integers had a sum of 96, what would be the largest integer?
Please show how to do it!!

Answer by Alan3354(69443)

If four consecutive odd integers had a sum of 96, what would be the largest integer?
-----------
The hard way:
n + n+2 + n+4 + n+6 = 96
4n+12 = 96
n = 21
===========
Better way:
96/4 = 24, the center.
2 below & 2 above
--> 21, 23, 25 & 27